OpenStudy (anonymous):

HELP PLZ ! If the r=8 units and h=4 units what is the voulme of the cylinder shown above using the formula. A.602.88 cubic units B.1,607.68 cubic units C.401.92 cubic units D.803.84 cubic units

OpenStudy (ybarrap):

 $$ v=2\pi r\times h $$

OpenStudy (sydthekid913):

the formula for the volume of a cylinder is pi * r^2 * h so if we replace the variables with the values given the formula would be pi * 8^2 * 4

OpenStudy (anonymous):

so the answer is B ? yall

OpenStudy (ybarrap):

Correction $$ V=\pi r^2\times h\\ =\pi 8^2 \times 4=128\pi $$

OpenStudy (anonymous):

@FateRocks8D

OpenStudy (anonymous):

@SydtheKid913 do u kno the answer in a rush plz

OpenStudy (anonymous):

o_o da sh%t I haven't done this in class sorry baby girl

OpenStudy (anonymous):

its okay thanks doe ;)

OpenStudy (sydthekid913):

ya ill help :P

OpenStudy (sydthekid913):

8 squared equals 64 64 time 4 equals 256 256 times pi equals 803.84 there you go your answer is D medal pwease?